II – PROBLEM SOLVING:(week 7)

1. The average public high school has 468

students with a standard deviation of 

87.

a) If a public school is selected, what 

is the probability that the number of 

students enrolled is greater than 

400?

b) If a random sample of 38 public 

elementary schools is selected, what 

is the probability that the number of 

students enrolled is between 445

and 485?



1
Expert's answer
2021-06-21T10:47:55-0400
  1. "\mu=468,\sigma=87"


(a) "P(X>400)=P(Z>\dfrac{400-468}{87})=P(z>-0.7816)=0.782"


(b)n=30,


"P(445<X<485)=P(\dfrac{445-468}{\frac{87}{\sqrt{38}}}<Z<\dfrac{485-468}{\frac{87}{\sqrt{38}}})\\[9pt]=P(-1.608<Z<1.188)=0.828"